OpenStudy (anonymous):

please help to solve 2/x+1+2x-3/x=0

OpenStudy (saifoo.khan):

\[\frac{2}{x+1} +\frac{2x-3}{x} = 0\]is it like this?

OpenStudy (anonymous):

yes sorry not good at typing

OpenStudy (hba):

What about taking lcm ?

OpenStudy (anonymous):

When working a problem like this, you get rid of the fractions. Do you know how to do that?

OpenStudy (anonymous):

not 100%

OpenStudy (anonymous):

Well try it and we'll guide you

OpenStudy (anonymous):

ok

OpenStudy (hba):

start solving it and we'll help you where you go wrong :)

OpenStudy (anonymous):

ok thanks

OpenStudy (shamim):

i wanna solve it

OpenStudy (saifoo.khan):

Are you still stuck? @strats1971

OpenStudy (anonymous):

yes i havnt a clue to be honest. cant work it out

OpenStudy (shamim):

\[2/(x+1)+(2x-3)/x=0\]

OpenStudy (saifoo.khan):